OpenStudy (anonymous):

help plse !anybody can answer this question One black coloured ball is surrounded by similar balls of red colour and again green colour. This pattern is repeated 10 times. Find the difference between total number of green and red balls.

OpenStudy (anonymous):

no

OpenStudy (anonymous):

A classroom has 26 rows (From A to Z), each row is divided in three columns and each column has five seats. To enter in the class, four doors (D1, D2, D3 and D4) are available and reaching up to the door two stairs (LH & RH) are available. A student wants to take the seat in the row C, F or P and then in how many ways he can get the seat.

OpenStudy (anonymous):

If, (P- 1)/9 = r, where r is triangular no and can be maximum of four digit no. What is the probability that P is a perfect number?

OpenStudy (anonymous):

Are these from statistics?

OpenStudy (anonymous):

I wouldn't know where to begin

OpenStudy (anonymous):

One black coloured ball is surrounded by similar balls of red colour and again green colour. This pattern is repeated 10 times. Find the difference between total number of green and red balls.

OpenStudy (anonymous):

I wouldn't know where to begin.

OpenStudy (anonymous):

one black is the first question

OpenStudy (anonymous):

I've not done these before

OpenStudy (anonymous):

A classroom has 26 rows (From A to Z), each row is divided in three columns and each column has five seats. To enter in the class, four doors (D1, D2, D3 and D4) are available and reaching up to the door two stairs (LH & RH) are available. A student wants to take the seat in the row C, F or P and then in how many ways he can get the seat.

OpenStudy (anonymous):

this one

OpenStudy (anonymous):

26*15=390,15/390+15/390+15/390,45/390,9/78 is it right

OpenStudy (compassionate):

What exactly are you having trouble with?

OpenStudy (anonymous):

A classroom has 26 rows (From A to Z), each row is divided in three columns and each column has five seats. To enter in the class, four doors (D1, D2, D3 and D4) are available and reaching up to the door two stairs (LH & RH) are available. A student wants to take the seat in the row C, F or P and then in how many ways he can get the seat.

OpenStudy (anonymous):

45/390 right?

OpenStudy (anonymous):

if u cannot try this problem atleast tis one.If, (P- 1)/9 = r, where r is triangular no and can be maximum of four digit no. What is the probability that P is a perfect number?

OpenStudy (compassionate):

@phi @ganeshie8 @Nurali @shamil98 It is 2:30 AM here. I am very tired.

OpenStudy (anonymous):

whenever possible u can answer.k u take rest

OpenStudy (anonymous):

@ganeshie8 is not responding

ganeshie8 (ganeshie8):

If, (P- 1)/9 = r, where r is triangular no and can be maximum of four digit no. What is the probability that P is a perfect number? wat are the perfect numbers less than 10000 ?

OpenStudy (anonymous):

6,28,496,8128

OpenStudy (anonymous):

i think 55 is the anwser

ganeshie8 (ganeshie8):

probability is always a number between 0 and 1

OpenStudy (anonymous):

(55-1)/9=6 is a triangular number

ganeshie8 (ganeshie8):

from where u got 55 ?

OpenStudy (anonymous):

sry i thought that r is a triangular no so 6 is a triangle no.so i told

ganeshie8 (ganeshie8):

(P- 1)/9 = r here, r is a triangular number, and P is a perfect number.

ganeshie8 (ganeshie8):

plug in ur perfect numbers one by one and see how many give u a triangular number

ganeshie8 (ganeshie8):

Perrfect numbers less than 10000 : 6, 28, 496, 8128 clearly 6 will not work. lets test 28 :- (28-1)/9 = 27/9 = 3 YES, 3 is a triangular number !

ganeshie8 (ganeshie8):

can u test remaning two ?

OpenStudy (anonymous):

496 is also perfect no but i dont know whether it is a triangular no

ganeshie8 (ganeshie8):

we dont care whether 496 is triangular or not

ganeshie8 (ganeshie8):

we need to check if (496-1)/9 is triangular or not. ok ?

ganeshie8 (ganeshie8):

lets test 496 :- (496-1)/9 = 495/9 = 55 = 10*11/2 = triangular number !!

OpenStudy (anonymous):

ya 55 is a triangular number

ganeshie8 (ganeshie8):

next, test 8128 :- (8128-1)/9 = 8127/9 = 903 = 42*43/2 = triangular number !

ganeshie8 (ganeshie8):

so, 28, 496 and 8128 satisfy the given relation that (P-1)/9 = r such that r is a triangular number

OpenStudy (anonymous):

yes it satisfies

ganeshie8 (ganeshie8):

next, how many triangular numbers are there less than 10000 ?

ganeshie8 (ganeshie8):

to find that, you need to solve below :- n(n+1)/2 < 10000 solving gives n <= 140

ganeshie8 (ganeshie8):

so probability = 3/140

ganeshie8 (ganeshie8):

see if that makes some sense

OpenStudy (anonymous):

yes right

OpenStudy (anonymous):

for this question above

OpenStudy (anonymous):

A classroom has 26 rows (From A to Z), each row is divided in three columns and each column has five seats. To enter in the class, four doors (D1, D2, D3 and D4) are available and reaching up to the door two stairs (LH & RH) are available. A student wants to take the seat in the row C, F or P and then in how many ways he can get the seat.

ganeshie8 (ganeshie8):

here is small trick to get ur questions answered quick by others :- ask oly ONE question per post :) close the questiona nd start a new thread so that, others get to see and assist u....

ganeshie8 (ganeshie8):

I had a look, its bit involved... il work it later ok

ganeshie8 (ganeshie8):

post it as new quesiton, lets see if somebody answers.... :)

OpenStudy (anonymous):

but no one answered wen i asked hte first q .so i posted many